OpenStudy (anonymous):

Write a rule for the linear function in the table. x f(x) –3 –1 0 2 3 5 6 8

OpenStudy (jlastino):

This is is similar form the problem before take two points (-3,-1) and (0,2) Use the two point formula y+1=(2+1)/(0+3) (x+3) y+1= 3/3 (x+3) y+1= 1(x+3) y+1= x+3 y=x+3-1 y=x+2 f(x)=x+2

OpenStudy (anonymous):

thanks